# Survey of Computerized Adaptive Testing: A Machine Learning Perspective

### 05 Apr 2024 v2

@@ 3.2. Categorization
As shown in Figure  3, we categorize existing CAT work into four parts based on the above three main components and the test control in CAT applications: 1) Cognitive Diagnostic Model, 2) Selection Algorithm, 3) Question Bank Construction, and 4) Test Control. Each part is further divided based on the different techniques employed. The following four sections (Section 4-7) will provide separate introductions and reviews of the relevant research within these four categories.
4. Cognitive Diagnosis Model
@@
The existing methods for CDM can be categorized into three main types: Latent Trait Model, Diagnostic Classification Model, and Deep Learning Model. Such classification is grounded on the nature of proficiency representation ($\theta$) in CAT—ranging from an overall numerical ability value (Latent Trait Model), to discrete cognitive states across different knowledge concepts (Diagnostic Classification Model), to a unified modeling approach via deep learning techniques (Deep Learning Model). The choice of model should depend on the specific goals of the assessment, the nature of the data, and the resources available. Regardless of the chosen CDM for estimating proficiency, the objective remains consistent: to minimize the error between the estimate and the true value at each step, expressed as $\|{\hat{\theta}^{t}}-\theta_{0}\|\to 0$.
4.1. Latent Trait Model
@@
In latent trait models, the examinee’s proficiency is handled as overall ability. Item Response Theory (IRT) (Embretson and Reise, 2013) is a representative work. In IRT, the examinee’s proficiency is modeled as scalars, which can be denoted as $\theta$. One of the most widely used models in IRT is the Three-Parameter Logistic Model (3PL-IRT). It utilizes a logistic-like interaction function to model the probability of examinee’s correct response to question $j$, i.e.,
(4) $f(q_{j},\theta)=c_{j}+\frac{1-c_{j}}{1+e^{-\alpha_{j}(\theta-\beta_{j})}}.$
@@
The 3PL-IRT model introduces three parameters ($\beta_{j},\alpha_{j},c_{j}$) for each test question $j$: The difficulty parameter $\beta_{j}$ corresponds to the level of proficiency at which an examinee has a 50% chance of answering the question correctly; The discrimination parameter $\alpha_{j}$ describes how well the question differentiates between examinees with different ability; The guessing parameter $c_{j}$ represents the probability that an examinee with a very low proficiency will answer the question correctly. In CAT systems, these parameters are pre-calibrated and remain fixed during the testing process, with annotation and calibration methods detailed in Section 6. IRT takes into account the number of questions answered correctly and the difficulty of the question. Today, almost all major adaptive tests, such as SAT and GRE, are developed by using IRT, because the methodology can significantly improve measurement reliability and interpretability (An and Yung, 2014). Multidimensional IRT (MIRT) (Reckase, 2006), on the other hand, extends IRT to multiple dimensions, allowing for the modeling of multiple latent traits simultaneously. Despite the great interpretability of (M)IRT models, their performance is constrained by the simplicity of the interaction function, and they lack fine-grained modeling about examinee’s cognitive states on individual knowledge concepts.
4.2. Diagnostic Classification Model
@@
Diagnostic Classification Model (DCM) is another representative class of CDMs, focusing on discrete knowledge concepts. Specifically, in DCMs, examinee proficiency is knowledge concept-wise and usually dichotomous, which indicates whether an examinee has mastered a knowledge concept or not. The DINA method (De La Torre, 2009; Von Davier, 2014) models examinee proficiency $\theta=\{\theta_{(1)},\theta_{(2)},...,\theta_{(K)}\}$ as their dichotomous knowledge mastery levels on all $K$ concepts. Given the Q-matrix $Q\in\mathbb{R}^{|\mathcal{Q}|\times K}$which is a binary matrix that indicates which knowledge concepts are associated with a question in bank $\mathcal{Q}$. DINA method focuses only on the knowledge concepts related to the target question $j$, where $Q_{jk}=1$. Thus, the examinee’s binary response variable (with proficiency $\theta$) to question $j$ is $\prod_{k,Q_{jk}=1}\theta_{(k)}$, and models questions as “slip” and “guess” parameters:
(5) $f(q_{j},\theta)=(1-s_{j})^{\prod_{k,Q_{jk}=1}\theta_{(k)}}g_{j}^{1-\prod_{k,Q_{jk}=1}\theta_{(k)}},$
@@
where $s_{j}$ is the slip parameter, indicating the likelihood of an incorrect response despite mastery, and $g_{j}$ is the guess parameter, reflecting the chance of a correct guess in the absence of mastery. Its extension G-DINA(de la Torre, 2011) provides a granular view of examinee proficiency, while FuzzyCDF(Liu et al., 2018) leverages fuzzy set theory for nuanced diagnostics from both objective and subjective data. Another approach, the Attribute Hierarchy Method (AHM)(Leighton et al., 2004), applies rule space theory to structure knowledge dependencies and align examinee proficiencies with the nearest ideal cognitive patterns to obtain diagnostic results. Compared to Latent Trait Models, Diagnostic Classification Models offer a more granular and comprehensive assessment of examinee proficiencies. They are particularly adept at providing detailed feedback on an individual’s strengths and weaknesses across multiple knowledge concepts, which is important for CAT and targeted further interventions. These models underscore a critical shift towards a more nuanced understanding of learning and proficiency, recognizing the multifaceted nature of knowledge acquisition and adaptive testing.
4.3. Deep Learning Model
@@
In recent years, the rapid growth of deep learning techniques stimulates the development of deep learning-driven CDMs. Compared to traditional models, deep learning methods are more suitable for cognitive diagnosis in large-scale data scenarios (e.g., online learning platforms) due to their efficiency and ability to learn the complex interaction pattern between examinees and questions. For instance, DIRT (Cheng et al., 2019) uses a neural network to capture semantic information from question texts to empower accuracy. NeuralCD (Wang et al., 2020c) utilizes a non-negative full connection neural network to capture the complex interaction, with the ability to generalize to other CDM methods. Considering the complex heterogeneous relationships between examinees, questions, and knowledge concepts, massive efforts have also been made to leverage them to enhance cognitive diagnosis (Gao et al., 2021; Li et al., 2022; Gao et al., 2022).
Discussion 1.
@@
In the CAT process, only a limited number of examinee responses can be obtained for proficiency estimation, due to the adaptive nature of the test. To some extent, CAT can be viewed as a cognitive diagnosis under a cold start scenario. The performance of the CDM is a critical factor in ensuring the accuracy of proficiency estimations within CAT. Meanwhile, it is important to note that the choice of the CDM can significantly influence the selection of corresponding question selection algorithm.
5. Selection Algorithm
@@
The selection algorithm is CAT’s core of implementing adaptivity and is the focal point of this survey. It utilizes the proficiency estimate obtained from the CDM (introduced in the above section) to choose the next most suitable question, ensuring an accurate estimation of proficiency while using the fewest possible questions. Question selection algorithms can be categorized into traditional methods based on statistical information, as well as more recent machine learning methods, e.g., data-driven approaches (i.e., Reinforcement Learning and Meta Learning), and Subset Selection are becoming increasingly prevalent.
5.1. Statistical Algorithms
In this subsection, we first establish the general framework and methodology for statistical question selection, which typically involves defining a specific information metric, e.g., Fisher information, or Kullback-Leibler information.
5.1.1. General Framework
@@
In general, a practical approach to designing a selection algorithm involves developing quantitative methods to assign a numerical value to each question in the bank $\mathcal{Q}$. The computer can then use this value assignment as a heuristic function to continuously select the most valuable questions during the personalized testing process, achieving adaptive question selection. In a more systematic view, well-adopted by classical question selection strategies, is from a statistical perspective, defining the value of a question as the informativeness it provides about the examinee’s potential ability estimation. The next question index $j_{t+1}$ can be selected from bank $\mathcal{Q}$ based on current estimate $\hat{\theta}^{t}$:
(6) $j_{t+1}=\arg\max_{q_{j}\in\mathcal{Q}}\mathcal{I}_{j}(\hat{\theta}^{t}),$
@@
where $\mathcal{I}_{j}(\cdot)$ is the informativeness of question $q_{j}$ (e.g., Fisher information). As illustrate in Definition 1, CAT assumes that each examinee has a true proficiency value ($\theta_{0}$) and it is considered as a parameter estimation process. The informativeness of a question can thus be interpreted as the expected contribution of the response on this question to the parameter estimation. This concept will be reflected in various selection algorithms discussed later. Fisher Information. In the parameter estimation problems, Fisher Information (Rissanen, 1996) is a concept from information theory and statistics that measures the amount of information that an observable random variable carries about an unknown parameter upon which the probability of that random variable depends. In the context of CAT, Fisher Information is often used to quantify the amount of information that a question provides about an examinee’s proficiency (Cheng, 2008). Specifically, we consider a random variable $X_{j}=(q_{j},y_{j})$ for which the pdf or pmf is $f(q_{j},\theta)$, where $\theta$ is the unknown parameter. The fisher info contained in the random variable $X_{j}$ is defined as: $\mathcal{I}_{j}(\theta)=\mathbb{E}_{y_{j}}[(\nabla_{\theta}L(X_{j}|\theta))^{2}]=\frac{(\nabla_{\theta}f(q_{j},\theta))^{2}}{f(q_{j},\theta)(1-f(q_{j},\theta))}$, where $L(X|\theta)={y\log f(q,\theta)+(1-y)\log(1-f(q,\theta))}$ is the likelihood function of $X$ with respect to the parameter $\theta$. Thus, when using 3PL-IRT to model the $f(q,\theta)$ and given current estimate $\hat{\theta}^{t}$, the Fisher Information of question $j$ can be calculated as:
@@
One crucial property of Fisher information is that its reciprocal (matrix inverse), is the variance (covariance matrix) of the asymptotic distribution of the maximum likelihood estimate:
Theorem 1 (The asymptotic distribution of MLE proficiency estimate (Ross, 2014)).
@@
Obviously, it can be obtained that as the number of test questions ($t$) or the Fisher information ($\mathcal{I}$) increases, the variance decreases. Since the estimated value is asymptotically unbiased (i.e., its mean equals the true value $\theta_{0}$), a decrease in its variance results in the distribution becoming more “tight” (higher $p(\hat{\theta}^{t}=\theta_{0})$), thereby reducing the uncertainty of the estimated proficiency $\hat{\theta}^{t}$. Due to the unknown $\theta_{0}$, maximizing the Fisher Information (MFI) at the current estimate $\hat{\theta}^{t}$ essentially serves to lower the variance of the proficiency estimate and reduce uncertainty, thereby improving the estimation efficiency. MFI has been widely popular in the development of personalized testing over the decades and extensively applied in practice as well. However, the effectiveness of MFI is most pronounced when the estimated value is close to the unknown true value $\theta_{0}$ (Chang, 2015). Therefore, using the current estimate in place of $\theta_{0}$ can sometimes lead to inaccuracies if the current estimate is far from the true value. Fisher Information is a mathematical concept that exists independently of CAT. Notably, this concept finds a natural extension in Multidimensional IRT (MIRT), where it transitions from a real number to a matrix (Hooker et al., 2009), facilitating the assessment of multiple abilities or traits simultaneously. Specifically, the information matrix provided by question $j$ at proficiency $\theta$ (now a vector) is defined as: $\mathcal{I}_{j}(\theta)=\mathbb{E}_{y_{j}}\left[\nabla\log L(X_{j}|\theta)\nabla\log L(X_{j}|\theta)^{T}\right]$. This Fisher Information matrix is instrumental in MIRT-based personalized assessment systems, as its inverse corresponds to the covariance matrix for maximum likelihood proficiency estimates. Researchers have exploited this relationship to develop many question selection strategies such as D-Optimization, A-Optimization, and E-Optimization, which respectively focus on maximizing the matrix’s determinant, minimizing the trace of its inverse, and maximizing its minimum eigenvalue (Reckase, 2006). These strategies, complemented by Bayesian optimization variants, have considerably advanced the statistical techniques available for question selection in adaptive testing. Kullback-Leibler Information. Fisher information is a commonly used measure in CAT due to its theoretical foundation, simplicity in mathematical form, and computational efficiency. However, the application of Fisher information might not be as efficient as expected if the estimators are not close to the true value $\theta_{0}$ (Chang, 2015). This issue becomes particularly evident during test’s initial stage, when the test length (number of questions) is too short to provide an accurate proficiency estimate. Consequently, relying on the MSI may not yield a satisfactory results, because Fisher information, often referred to as local information, only considers the information at a specific point ($\hat{\theta}^{t}$). To address this issue, Chang and Ying (1996) proposed a global information measure based on Kullback-Leibler (KL) divergence, also known as KL information: Let $\theta_{0}$ be the true proficiency, for any $\theta$, the KL information of the question $q_{j}$ (with response $X_{j}$) which measures the discrepancy between the two probability distributions specified by $\theta_{0}$ and $\theta$, is defined by
$KL_{j}(\theta||{\theta}_{0})=\mathbb{E}_{y_{j}}\log\left[\frac{L(X_{j}|\theta_{0})}{L(X_{j}|\theta)}\right]=f(q_{j},\theta_{0})\log\left[\frac{f(q_{j},\theta_{0})}{f(q_{j},\theta)}\right]+(1-f(q_{j},\theta_{0}))\log\left[\frac{1-f(q_{j},\theta_{0})}{1-f(q_{j},\theta)}\right].$
The important feature is that KL is a function of two levels, $\theta$ and $\theta_{0}$. $KL(\theta||{\theta}_{0})$ represents the discrimination power of the question on the two levels and does not require that $\theta$ be close to $\theta_{0}$. The corresponding algorithm for question selection that maximizes KL information involves integrating over $\theta$ in Eq.(5.1.1) and substituting the current estimate for $\theta_{0}$:
(7) $\mathcal{I}_{j}(\hat{\theta}^{t})=\int_{\hat{\theta}^{t}-\delta}^{\hat{\theta}^{t}+\delta}KL_{j}(\theta||\hat{\theta}^{t})d\theta.$
where $\delta=3/\sqrt{t}$ and the integration range is relatively large at the test’s outset and gradually narrows as $t$ increases. In MIRT, the global KL divergence information extends from a univariate integral to a multivariate integral.
Figure 4. Illustration of KL and Fisher information functions for two questions (Question 1: $\alpha=1.7,\beta=1.9,c=0.1$; Question 2: $\alpha=2.3,\beta=0.5,c=0.3$). Assuming the current proficiency estimate $\hat{\theta}^{t}=1$. The KL information (left) for the given question represents an integral centered around $\hat{\theta}^{t}$, while the Fisher information (right) corresponds to the value at the specific point $\hat{\theta}^{t}$.
Essentially, the KL method identifies questions that will provide the greatest differentiation between the examinee’s possible proficiency levels. Figure 4 illustrates the KL and Fisher information functions for two distinct questions. The information conveyed by KL Information is not confined to a single point value like Fisher; instead, it spans a wide range of proficiency levels $\theta$. This is why KL information is global while Fisher information is local (Chang, 2015). Interestingly, for $\theta$ near $\theta_{0}$, KL Information and Fisher information are always close. If we envision KL Information as a curve, Fisher information corresponds to its curvature (second derivative) at $\theta=\theta_{0}$. This suggests that Fisher information can be derived from KL Information, but the converse is not true.
5.1.2. Advances and Generalization
[…]
Discussion 2.
@@ 5.1.2. Advances and Generalization
Selection algorithms in CAT have primarily relied on the above statistical heuristic approaches, which require domain experts to consider every possible testing scenario and manually design corresponding selection algorithms. These methods are model-specific, requiring distinct selection strategies for different CDMs. For example, the above MFI strategy (Lord, 2012) is specifically crafted for (M)IRT. Consequently, previous statistical methods lack flexibility and are labor-intensive, and the selection algorithm must be re-designed if the underlying CDM changes.
5.2. Active Learning Algorithms
An ideal CAT framework should be model-agnostic, meaning that the selection algorithms can adapt to different CDMs while remaining effective. Researchers are exploring a broader approach to the CAT problem by investigating its correlation with a versatile machine learning technique—Active Learning (Krishnakumar, 2007). Active Learning is to choose valuable data actively, thus can train better models with less data. This method has improved data efficiency and cut annotation costs in numerous learning tasks (Budd et al., 2021). As shown in Figure 5, in active learning, each cycle begins with a learning model and a strategy for selecting data. The system selects an unlabeled data (set) to label at each iteration, using this to supplement the limited labeled data and improve the model’s performance. The key point behind this is how to design the data selection strategy. Existing methods typically view selection from two perspectives: informativeness and representativeness. Informativeness-based algorithms choose samples that significantly reduce a learning model’s uncertainty (Yoo and Kweon, 2019). Representativeness-based algorithms, on the other hand, choose samples that best reflect the overall patterns of unlabeled data (Ghorbani et al., 2022; Wang et al., 2020b), and their combinations (Wang et al., 2020b).

### 09 Mar 2026 v3

@@ 3.2 Categorization
As shown in Figure 2, we categorize existing CAT research into four major components involved in the testing process described above: 1) Measurement Model, 2) Selection Algorithm, 3) Question Bank Construction, and 4) Test Control. Each part is further divided based on the different techniques employed. The following four sections (Sections 47) provide detailed introductions and literature reviews of these components, with a particular focus on their theoretical foundations and methodological developments from a machine learning perspective.
4 Measurement Model
The existing methods for Measurement Model can be categorized into three main types: Item Response Theory (IRT), Cognitive Diagnostic Model (CDM), and Deep Learning Model. In the first 50 years of CAT development, IRT was the dominant modeling framework and widely adopted in operational systems. It was not until 2009, with the introduction of CD-CAT [32], that CDM began to be used as the underlying measurement model in adaptive testing. More recently, with the increasing scale and complexity of adaptive assessments and the rise of deep learning, a variety of new measurement models have emerged that go beyond traditional IRT and CDM.
@@ 4 Measurement Model
Such classification is grounded on the nature of proficiency representation ($\theta$) in CAT—ranging from an overall numerical ability value (i.e., IRT), to discrete cognitive states across different knowledge concepts (i.e., CDM), to a unified modeling approach via deep learning techniques (i.e., Deep Learning Model). The choice of model should depend on the specific goals of the assessment, the nature of the data, and the resources available. Regardless of the chosen Measurement Model for estimating proficiency, the objective remains consistent: to minimize the error between the estimate and the true value at each step, expressed as $\|{\hat{\theta}^{t}}-\theta_{0}\|\to 0$.
4.1 Item Response Theory
@@
In IRT [49], an examinee’s proficiency is typically represented as a continuous scalar variable, referred to as overall ability. As a foundational framework in measurement models, IRT represents the examinee’s general level of proficiency using a latent trait parameter $\theta$. One of the most widely used models in IRT is the Three-Parameter Logistic Model (3PL-IRT). It utilizes a logistic-like interaction function to model the probability of examinee’s correct response to question $j$, i.e.,
$f(q_{j},\theta)=c_{j}+\frac{1-c_{j}}{1+e^{-\alpha_{j}(\theta-\beta_{j})}}.$ (4)
@@
The 3PL-IRT model introduces three parameters ($\beta_{j},\alpha_{j},c_{j}$) for each test question $j$: The difficulty parameter $\beta_{j}$ corresponds to the level of proficiency at which an examinee has a 50% chance of answering the question correctly; The discrimination parameter $\alpha_{j}$ describes how well the question differentiates between examinees with different ability; The guessing parameter $c_{j}$ represents the probability that an examinee with a very low proficiency will answer the question correctly. In CAT systems, these parameters are pre-calibrated and remain fixed during the testing process, with annotation and calibration methods detailed in Section 6. IRT takes into account the number of questions answered correctly and the difficulty of the question. Almost all major adaptive tests for humans, such as SAT and GRE, are developed by using IRT, because the methodology can significantly improve measurement reliability and interpretability [3]. Recently, for AI system evaluation, Polo et al. [125] successfully selected 100 informative curated questions from MMLU [67], a popular multiple-choice QA benchmark consisting of 14K questions, and accurately estimated the performance of LLMs.
Multidimensional IRT (MIRT) [127], on the other hand, extends IRT to multiple dimensions, allowing for the modeling of multiple latent traits simultaneously. Despite the great interpretability of (M)IRT models, their performance is constrained by the simplicity of the interaction function, and they lack fine-grained modeling about examinee’s cognitive states on individual knowledge concepts.
4.2 Cognitive Diagnostic Model
@@
Cognitive Diagnostic Model (CDM) is another representative class of measurement models, focusing on discrete knowledge concepts. Specifically, in CDMs, examinee proficiency is knowledge concept-wise and usually dichotomous, which indicates whether an examinee has mastered a knowledge concept or not. For example, in a mathematics assessment, knowledge concepts may include addition, fractions, or solving linear equations. We continue to use $\theta$ to denote examinee proficiency for consistency. For example, the DINA method [40, 164] models examinee proficiency $\theta=\{\theta_{(1)},\theta_{(2)},...,\theta_{(K)}\}$ as their dichotomous knowledge mastery levels on all $K$ concepts. Given the Q-matrix $Q\in\mathbb{R}^{|\mathcal{Q}|\times K}$which is a binary matrix that indicates which knowledge concepts are associated with a question in bank $\mathcal{Q}$. DINA method focuses only on the knowledge concepts related to the target question $j$, where $Q_{jk}=1$. Thus, the examinee’s binary response variable (with proficiency $\theta$) to question $j$ is $\prod_{k,Q_{jk}=1}\theta_{(k)}$, and models questions as “slip” and “guess” parameters:
$f(q_{j},\theta)=(1-s_{j})^{\prod_{k,Q_{jk}=1}\theta_{(k)}}g_{j}^{1-\prod_{k,Q_{jk}=1}\theta_{(k)}},$ (5)
@@
where $s_{j}$ is the slip parameter, indicating the likelihood of an incorrect response despite mastery, and $g_{j}$ is the guess parameter, reflecting the chance of a correct guess in the absence of mastery. Its extension G-DINA[41] provides a granular view of examinee proficiency, while FuzzyCDF[102] leverages fuzzy set theory for nuanced diagnostics from both objective and subjective data. Another approach, the Attribute Hierarchy Method [89], applies rule space theory to structure knowledge dependencies and align examinee proficiencies with the nearest ideal cognitive patterns to obtain diagnostic results.
Compared to IRT, CDM offers a more granular and comprehensive assessment of examinee proficiencies. They are particularly adept at providing detailed feedback on an individual’s strengths and weaknesses across multiple knowledge concepts, which is important for CAT and targeted further interventions. These models underscore a critical shift towards a more nuanced understanding of learning and proficiency, recognizing the multifaceted nature of knowledge acquisition and adaptive testing.
4.3 Deep Learning Model
@@
In recent years, the rapid growth of deep learning techniques stimulates the development of deep learning-driven Measurement Models. Compared to traditional models, deep learning methods are more suitable for measurements in large-scale data scenarios (e.g., online learning platforms) due to their efficiency and ability to learn the complex interaction pattern between examinees and questions.
In these models, an examinee’s proficiency $\theta$ is typically represented by a high-dimensional latent vector (embedding). Similarly, each question is encoded as an question embedding $e_{j}=\text{Embed}(q_{j})$. These embeddings are passed through a multi-layer neural network to predict the probability of a correct response:
$f(q_{j},\theta)=\phi_{n}\left(\cdots\phi_{1}\left(W[\theta;e_{j}]+b\right)\cdots\right),$ (6)
where $W$ and $b$ are the weight matrix and bias vector, respectively, and $\phi_{k}(\cdot)$ denotes the activation function at the $k$-th layer (e.g., ReLU, Tanh, or Sigmoid).
Based on this framework, several deep learning-based measurement models have demonstrated strong performance. For example, DIRT [30] uses a neural network to capture semantic information from question texts to empower accuracy. NeuralCD [167] utilizes a non-negative full connection neural network to capture the complex interaction, with the ability to generalize to other measurement models. Considering the complex heterogeneous relationships between examinees, questions, and knowledge concepts, massive efforts have also been made to leverage them to enhance measurements [55, 92, 54].
@@
Discussion: In the CAT process, only a limited number of examinee responses can be obtained for proficiency estimation. To some extent, CAT can be viewed as a proficiency measurement under a cold start scenario. The performance of the measurement model is a critical factor in ensuring the accuracy of proficiency estimations within CAT. Meanwhile, it is important to note that the choice of the measurement model can significantly influence the selection of corresponding question selection algorithm.
5 Selection Algorithm
@@
The selection algorithm is CAT’s core of implementing adaptivity and is the focal point of this survey. It utilizes the proficiency estimate obtained from the Measurement Model (introduced in the above section) to choose the next most suitable question, ensuring an accurate estimation of proficiency while using the fewest possible questions. Question selection algorithms can be categorized into traditional methods based on statistical information, as well as more recent machine learning methods, e.g., data-driven approaches (i.e., Reinforcement Learning and Meta Learning), and Subset Selection are becoming increasingly prevalent.
5.1 Statistical Algorithms
@@
Generally, a practical approach to designing a selection algorithm involves developing quantitative methods to assign a numerical value to each question in the bank $\mathcal{Q}$. Classical statistical selection algorithms define the value of a question as the informativeness it provides about the examinee’s potential ability estimation. The next question index $j_{t+1}$ can be selected from bank $\mathcal{Q}$ based on current estimate $\hat{\theta}^{t}$:
$j_{t+1}=\arg\max_{q_{j}\in\mathcal{Q}}\mathcal{I}_{j}(\hat{\theta}^{t}),$ (7)
@@
where $\mathcal{I}_{j}(\cdot)$ is the informativeness of question $q_{j}$ (e.g., Fisher information). As illustrated in Definition 1, CAT assumes that each examinee has a true proficiency value ($\theta_{0}$) and it is considered as a parameter estimation process. The informativeness of an question can thus be interpreted as the expected contribution of the response on this question to the parameter estimation. This concept will be reflected in various selection algorithms discussed later.
Fisher Information. In the parameter estimation problems, Fisher Information [130] is a concept from information theory and statistics that measures the amount of information that an observable random variable carries about the unknown parameter. In CAT, Fisher Information is often used to quantify the amount of information that a question provides about an examinee’s proficiency [31]. Specifically, we consider a random variable $\mathcal{D}_{j}=(q_{j},y_{j})$ for which the pdf or pmf is $f(q_{j},\theta)$, where $\theta$ is the unknown parameter. The fisher info contained in the variable $\mathcal{D}_{j}$ is defined as: $\mathcal{I}_{j}(\theta)=\mathbb{E}_{y_{j}}[(\nabla_{\theta}L(\mathcal{D}_{j}|\theta))^{2}]=\frac{(\nabla_{\theta}f(q_{j},\theta))^{2}}{f(q_{j},\theta)(1-f(q_{j},\theta))}$, where $L(\mathcal{D}|\theta)={y\log f(q,\theta)+(1-y)\log(1-f(q,\theta))}$ is the likelihood function of $\mathcal{D}$ with respect to the parameter $\theta$.
Thus, when using 3PL-IRT to model the $f(q,\theta)$ and given current estimate $\hat{\theta}^{t}$, the Fisher Information of question $j$ can be calculated as:
@@
One crucial property of Fisher information is that its reciprocal (matrix inverse), is the variance (covariance matrix) of the asymptotic distribution of the proficiency estimate:
Theorem 1 (The asymptotic distribution of MLE proficiency estimate [136])
@@
Obviously, as the number of questions $t$ or the Fisher Information $\mathcal{I}(\theta_{0})$ increases, the variance of the estimate decreases. Since $\hat{\theta}^{t}$ is asymptotically unbiased (i.e., $\mathbb{E}[\hat{\theta}^{t}]=\theta_{0}$), a lower variance implies a more concentrated distribution around $\theta_{0}$, thereby reducing estimation uncertainty and improving the estimation efficiency.
Fisher information has been popular in the development of personalized testing over the decades and extensively applied in various standardized human assessments. Similarly, for AI model evaluations, particularly for LLMs, the simple Fisher method allows for accurate performance estimation using only a small sample of test data. For example, Kipnis et al. [81] reduce six commonly used benchmarks to less than 3% of their original size while accurately estimating the performance of over 5,000 LLMs.
When the measurement model is MIRT, Fisher information naturally extends from a scalar to a matrix [71]. Specifically, the information matrix provided by question $j$ at proficiency $\theta$ (now a vector) is defined as: $\mathcal{I}_{j}(\theta)=\mathbb{E}_{y_{j}}\left[\nabla\log L(\mathcal{D}_{j}|\theta)\nabla\log L(\mathcal{D}_{j}|\theta)^{\top}\right]$. This matrix’s inverse approximates the covariance of the MLE proficiency estimate. Based on this, various selection algorithms have been proposed, including D-Optimality (maximizing the determinant), A-Optimality (minimizing the trace of the inverse), and E-Optimality (maximizing the smallest eigenvalue) [127].
Kullback-Leibler Information. Fisher information is widely used in CAT for its theoretical foundation and mathematical simplicity. However, its effectiveness diminishes when the proficiency estimate deviates from the true value $\theta_{0}$ [25]. This issue becomes particularly evident in the early stages of a test when the estimate is still unstable due to limited responses. To address this issue, Chang et al. [23] proposed a global information measure based on Kullback–Leibler (KL) divergence. For the given question $q_{j}$ (with response $\mathcal{D}_{j}$), the KL divergence between a candidate proficiency level $\theta$ and the true proficiency $\theta_{0}$ is defined as:
$KL_{j}(\theta\|{\theta}_{0})=\mathbb{E}_{y_{j}}\log\frac{L(\mathcal{D}_{j}|\theta_{0})}{L(\mathcal{D}_{j}|\theta)}$
$=$ $f(q_{j},\theta_{0})\log\frac{f(q_{j},\theta_{0})}{f(q_{j},\theta)}+(1-f(q_{j},\theta_{0}))\log\frac{1-f(q_{j},\theta_{0})}{1-f(q_{j},\theta)}.$
The corresponding question selection algorithm integrates KL over a neighborhood of the current estimate $\hat{\theta}^{t}$:
$\mathcal{I}_{j}(\hat{\theta}^{t})=\int_{\hat{\theta}^{t}-\delta}^{\hat{\theta}^{t}+\delta}KL_{j}(\theta||\hat{\theta}^{t})d\theta.$ (8)
where $\delta=3/\sqrt{t}$. The integration range is wide at the beginning of the test and and gradually narrows as $t$ increases. In MIRT, this extends naturally to a multivariate integral. Essentially, the KL information identifies questions that will provide the greatest differentiation between the examinee’s possible proficiency levels. Unlike Fisher information, which depends on a single point estimate, KL information measures the discrepancy between two proficiency levels, $\theta$ and $\theta_{0}$, and remains effective even when they differ significantly. This is why KL information is global while Fisher information is local [25]. Specific examples comparing the two can be found in the appendix.
Advanced Statistical Algorithms. Numerous works based on Fisher and KL information have been proposed. These methods try to introduce more information in selection to improve the efficiency of proficiency estimation. The Maximum Likelihood Weighted Information [160] weights the Fisher information by the likelihood function of the examinee’s current response results. Its rationale is similar to KL information and aims to improve the local limitations of Fisher information: selecting the one that maximizes the integral of the likelihood function times the Fisher $\mathcal{I}_{j}(\theta)$ over the proficiency level: $\mathcal{I}_{j}(\hat{\theta}^{t})=\int_{\hat{\theta}^{t}-\delta}^{\hat{\theta}^{t}+\delta}L(\mathcal{D}_{1:t-1}|\theta)\mathcal{I}_{j}(\theta)d\theta$, where $L(\mathcal{D}_{1:t-1}|\theta)=\sum_{j=1}^{t-1}L(\mathcal{D}_{j}|\theta)$ is the likelihood function of previous $t$ response. Furthermore, the Maximum Posterior Weighted Information [158, 10] further weights the Fisher and KL information with an additional posterior probability distribution $P(\theta|\mathcal{D}_{1:t-1})$: $\mathcal{I}_{j}(\hat{\theta}^{t})=\int_{\hat{\theta}^{t}-\delta}^{\hat{\theta}^{t}+\delta}P(\theta|\mathcal{D}_{1:t-1})L(\mathcal{D}_{1:t-1}|\theta)\mathcal{I}_{j}(\theta)d\theta$, where $\mathcal{I}(\theta)$ can be the Fisher Information or KL divergence. Maximum Expected Information [158] accounts for all possible outcomes $y_{t}$ and their impact on the updated proficiency estimate when weighting Fisher information. Lastly, the theta-Optimization (thOpt) process [11] selects questions by aligning the maximized information with the current proficiency estimate. Some non-parametric machine learning methods (e.g., decision trees) can also be explored, showing that a small number of questions can match or exceed traditional methods in accuracy, especially under high-dimensional and imbalanced data conditions [184].
The selection algorithms discussed earlier are based on IRT and do not directly apply to other measurement models. As noted in Section 4, measurement models like CDM represent proficiency $\theta$ as discrete states across different knowledge concepts. Adaptive selection algorithms under such models is known as Cognitive Diagnosis CAT (CD-CAT) [32]. While the selection principles remain similar (maximizing information from selected items), the information measures differ. In CD-CAT, techniques based on KL divergence [68] and Shannon entropy [151] are commonly used. Although continuous traits (as modeled in IRT) and discrete states (as modeled in DINA) describe different aspects of proficiency, they are complementary. This has led to the development of dual-objective CD-CAT methods [79, 183, 38] that aim to assess both simultaneously .
@@
Discussion: Selection algorithms in CAT have primarily relied on the above statistical heuristic approaches, which require domain experts to consider every possible testing scenario and manually design corresponding selection algorithms. These methods are model-specific, requiring distinct selection algorithms for different measurement models. For example, the above Fisher information [106] is specifically crafted for (M)IRT. Consequently, previous statistical methods lack flexibility, and the selection algorithm must be re-designed if the underlying measurement model changes.